# arXiv:2602.13749v3[cond-mat.str-el] 02 Mar 2026

### v3

@@ III Perturbative renormalization group results
For a fixed $g_{r}$ and an increasing $g_{i}$ with gradually enlarged non-Hermiticity, the system firstly flows to the strong-coupling phase for small $\theta$ and then flows to the weak-coupling phase for large $\theta$. Since the perturbative RG method is only expected to be valid and reliable for weak-coupling phase, it is intriguing to ask whether more exotic phases could occur for the strong or intermediate coupling regime. Notably, recent Bethe ansatz studies have clearly implied that there indeed emerges an extra phase in the intermediate region for intermediate $\theta$ [35], where additional degrees of freedom screen the local moment. This is termed as the Yu-Shiba-Rusinov (YSR) phase owing to its similarity with impurity states in superconductors [111, 112]. Since this emerging new phase is missed by the perturbative RG analysis, it is obvious that more careful non-perturbative approach is indispensable to achieve a better and comprehensive understanding of the non-Hermitian Kondo effect.
IV Bethe ansatz solution
@@
where $v_{f}$ denotes the Fermi velocity, $\psi_{k,\alpha}^{\dagger}(x),\psi_{k,\alpha}(x)$ are the creation and annihilation operator of conduction electrons (after the linearization of the spectrum around Fermi surface and the mapping to 1D model [113, 114, 115, 116, 117, 118, 119]) defined at coordinate $x$ with spin $\alpha$ and channel $k$. $J$ is the bare complex coupling. The corresponding Bethe ansatz equations is obtained as,
@@
The non-Hermitian Kondo model is characterized by two RG-invariant parameters [35], i.e., the Kondo energy scale $T_{K}$ and the non-Hermicity parameter $\alpha$, given by,
@@
where $\nu=\frac{N}{L}$ is the density of electron. For simplicity, we set $\widetilde{g}=1$ in the following calculations.
@@
where we have use the identity, $\int\text{sech}(\frac{\pi x}{n})dx=n$. The total spin $S=0$ implies that the impurity is fully screened.
@@
The spin excitations carrying spin 1/2 are termed as the spinons. Note that the spinon energy is complex, which consists of both a real part $E_{h}$ and a imaginary part $\Gamma_{h}$.
@@
which is a function of $\alpha$ and $\frac{E}{T_{K}}$. We then plot $\nu_{s}$ for different $\alpha$, as shown in Fig.2(a). When $0\leq\alpha\leq\pi/6$, the spinon DOS remains a pure Kondo behavior with a peak at $E=0$. When $\alpha\geq\pi/6$, the peak moves to $E/T_{K}=\sqrt{2\sin\alpha-1}$, until $\alpha$ approaches $\pi/2$. Meanwhile, the spinon DOS develops a delta peak, i.e., $\delta(E-T_{K})/2$ at $\alpha=\pi/2$. These results coherently indicate that the number of modes that screen the impurity decreases with $\alpha$, until reaching the case where the screening is contribued only by a single mode at $E/T_{K}=1$.
@@
The above Bethe ansatz solution is obtained for the general non-Hermitian multichannel Kondo model with isotropic Kondo exchange couplings, thus is not applicable for the $\mathcal{PT}$ symmetric case. For the $\mathcal{PT}$ symmetric case (with $J^{\prime}_{3}=J^{\prime}$, $J^{\prime}_{1}=J^{\prime\star}_{2}=J^{\prime}e^{i\theta}$), although we do not have access to the specific Bethe ansatz solution yet, we expect that only the contribution of the real part of impurity scattering could remain, since the imaginary part contributed by the conjugated $J^{\prime}_{1}$, $J^{\prime\star}_{2}$ pairs may cancel with each other, as will be implied by the NHNRG results below. Here, for the $\mathcal{PT}$ asymmetric case, we clearly see that the non-Hermitian effects can induce exotic phases absent in the Hermitian counterparts. Then, a urgent question to be answered is that how to obtain clear signatures of the emerging YSR phase, particularly regarding the thermodynamic quantities. We will further clarify this question in the next section using a non-Hermitian numerical renormalization group (NHNRG) method.
V non-Hermitian nrg approach
V.1 Numerical implementations
In this section, we will demonstrate the details in the implementation of nHNRG, focusing on its difference with convention NRG for Hermitian systems.
The conventional NRG method involves several steps [125]. The first step is the logarithmical discretization. One reformulates the original impurity model using an one-dimensional (1D) energy representation for the bath electrons under a band width cutoff $\pm D$. The conduction band can be discretized to a series of energy intervals $D_{n}$ with a scaling parameter $\Lambda$, i.e.,
@@ V.1 Numerical implementations
where $H_{N}$ is the first $N$ site Hamiltonian, and $h_{N+1}$ represents for short-range terms containing the $(N+1)$-th site. In each iteration, we truncate the energy spectrum and dilate the kept eigenenergies, which gradually yields the low-energy physics. The success of this strategy is guaranteed by the energy scale separation of nearest sites in the Wilson chain, owing to the logarithmical discretization. The iteration diagonalization can then be stopped as long as a stable low-energy fixed point is reached. Such a diagonalization process also reveals information about the fixed points, as well as thermodynamic properties with varying energy scales.
@@
Although the two set of wavefunctions are not orthogonal by themselves alone, they do form a biorthogonal basis that satisifes, [127]
@@
From above, one observes that the differences between Hermitian and non-Hermitian NRG lies in their their different ways of diagonalization. For non-Hermitian systems, we need to calculate both left and right eigenvectors for each iteration step, as clarified in Eq.(57). This generates more detailed issues regarding the non-unitary transformation, truncation, and the degeneracy, etc, which will be explained in the following.
@@
When $\mathcal{PT}$ symmetry is not spontaneously broken, the energy spectrum will be strictly real. In this case, the metric $\eta$ is not only invertible but also positive. Thus, the thermodynamic quantities defined in Eq.(74) is completely valid and is analogous to those in conventional Hermitian systems [131, 127]. Whereas, when the $\mathcal{PT}$ symmetry is spontaneously broken, the eigenenergies will either be real or take place as complex conjugate pairs. Thus, the partition function $Z_{N}^{\rm LR}$ still maintains to be real, since we have $e^{-\beta E}+e^{-\beta E*}=2e^{-\beta{\rm Re}(E)}{\rm cos[Im}(\beta E)]$. In this case, it should be noted that the Boltzmann weight and the partition function is not positive-definite in general. Although this is forbidden for closed Hermitian systems, this feature reflects the essence of the non-Hermiticity, which physically originates from open systems with loss or gain. Moreover, for the model and parameters studied in this work, we will show in the following that the partition function maintains positive for all energy scales, as indicated by Fig.6(b).
@@
Now we discuss more numerical details about solving the left and right eigenstates, which is complicated by degeneracies [132]. The degeneracies often arise from symmetries, either at the Hamiltonian level or the emergent ones at low-energies. The problem can be partially overcome by decomposing the Hamiltonian into subspaces of different irreducible representations. Then, one only need to solve certain reduced block matrices. Since the size of the blocks are much smaller than the full matrix, the computational cost is also reduced [133, 134]. Taking the $n=2$ NHMCK model as an example, one can utilize the ${\rm U}(1)_{\rm charge}\times{\rm SU}(2)_{\rm channel}$ symmetry. However, for symmetries emergent around the low-energy fixed points [135], the degeneracy cannot be avoided. For non-Hermitian matrices, most of numerical eigensolver packages do not support a biorthonormal output. One has to orthogonalize and normalize the eigenstates manually in each degenerate subspace. An open-source program has been recently proposed for non-Hermitian NRG calculation [132] for simple models such as the single impurity Anderson model. In fact, one can first solve the right eigenvectors with common eigensolvers, and then obtain the left eigenvectors by the inverse of the right one, i.e.,
@@
Truncation is another key procedure in NRG calculations. In Hermitian cases, the usual scheme is just to drop out excited states higher than a energy cutoff $E_{\rm k}$, keeping the lowest $N_{\rm k}$ states. For non-Hermitian cases, the eigenenergies are generally complex and the cutoff needs a careful benchmark. In Ref. [132], several truncation schemes are discussed, and a “LowRe” scheme is demonstrated to yield the most stable and accurate numerical results given the cutoff $N_{\rm k}$. This scheme sorts the eigenvalues primarily by their real parts $\textrm{Re}(E)$ and secondarily by their imaginary parts $\textrm{Im}(E)$. We adopt the same cutoff scheme in our calculations. However, beyond Ref. [132], we focus on evaluating the thermodynamic quantities defined by Eq.(74) with varying temperatures for both the $\mathcal{PT}$-asymmetric and symmetric cases.
V.2 Non-Hermtitian NRG results for the PT asymmetric model
We now solve the NHMCK model taking $n=2$ (two-channel) as an example. Compared to models with more channels, it is of minimal computation cost and is already sufficient to demonstrate the key numerical features. The channel-symmetry is assumed since the proposed setup in Sec. II ensures exact symmetry between different channels. For the $\mathcal{PT}$-asymmetric model, we consider a general complex Kondo exchange coupling, $J=J_{0}e^{i\theta}$ with a tunable $\theta$ that quantifies the strength of non-Hermiticity. For the $\mathcal{PT}$-symmetric model, we assume anisotropic Kondo coupling that takes the form of ${\bf J}=(J_{x},J_{y},J_{z})=J_{0}(e^{i\theta},e^{-i\theta},1)$. Such an anisotropic model can also be realized by tuning the tunneling coefficient in the multi-junction setup, as discussed earlier. For $\theta=0$, both cases reduce to the conventional Hermitian two-channel Kondo model. In this section, we show the NRG results for the $\mathcal{PT}$-asymmetric model, while those for the $\mathcal{PT}$-symmetric model will be presented in the next section.
Figure 3: The energy spectrum for the $\mathcal{PT}$-broken $n=2$ NHMCK model obtained by non-Hermitian NRG for different $\theta$. The upper ((a)(c)(e)) and lower ((b)(d)(f)) panel show the real and imaginary part, i.e., Re(E)(E) and Im(E)(E), respectively. The maximal absolute value of the imaginary parts are marked for each plot in (b)(d)(f). The first 50 eigenenergies are plot here. The black curves show the calculated Re(E)(E) corresponding to the Hermitian two-channel Kondo model with $\theta=0$.
@@ V.2 Non-Hermtitian NRG results for the PT asymmetric model
The energy spectrum of the $\mathcal{PT}$-asymmetric model is shown in Fig.3. Notably, the spectrum deviates from the Hermitian result (the dark curves) as $\theta$ increases. For small $\theta$ , the real part of the spectrum Re(E)(E) agrees well with that of the Hermitian case (Fig.3(a)), with only a negligible imaginary part (Fig.3(b)). For intermediate $\theta$, Re(E)(E) displays obvious deviations at relatively high energy scales (Fig.3(c)), displaying larger imaginary part in the same energy range (Fig.3(d)). However, the system still flows to the Hermitian fixed point at low energies. For large $\theta$, the spectrum is completely distinct from the Hermitian, and the low-energy fixed point is also altered. The absolute values of the imaginary part Im($E$) also increase with $\theta$, as marked in Fig.3(b),(d),(f).
@@
Thermodynamic quantities for the $\mathcal{PT}$-asymmetric model also display interesting non-Hermitian effects. We calculate the impurity entropy $S_{\rm imp}$ and the impurity specific heat $C_{\rm imp}$ as a function of temperatures. As shown in Fig.4(a), the model in fact exhibits three different regimes as $\theta$ increases. The first is the Kondo regime with weak non-Hermiticity ($\theta=0.10$) where $S_{\rm imp}$ monotonically evolves from ln2 to $\frac{1}{2}$ln2. This indicates the (over)screening process of the impurity with lowering temperature, similar to the Hermitian case. The second is the intermediate non-Hermticity regime ($\theta=0.50$), where $S_{\rm imp}$ also evolves from ln2 to $\frac{1}{2}$ln2, but clearly in a non-monotonous behavior. The third is the strong non-Hermticity regime ($\theta=1.00$), where the impurity entropy $S_{\rm imp}$ is close to ln 2 in both the UV and IR limit, displaying the local moment phase.
@@
We further compare the imaginary parts of the partition function $Z$ and the impurity entropy $S_{\rm imp}$ to their real parts in Fig.4(b) and (c). We find that the imaginary parts vanish as the NRG iteration goes to the IR fixed point, but can not be ignored at higher energies. Furthermore, the nonzero imaginary parts hallmark the characteristic energy scale where real parts of the impurity thermodynamic properties show YSR-like behaviors. The non-monotonic entropy and the overshoot of $\mathrm{ln}2$ indicate that there are extra DOF with finite energy contributing to the impurity thermodynamics, which are totally induced by the non-Hermitian effects.
Figure 5: The nHNRG results of the energy spectrum for the $\mathcal{PT}$-symmetric $n=2$ NHMCK model for different $\theta$.
VI the pseudo-Hermitian model with PT symmetry
VI.1 Non-Hermitian NRG results for the PT symmetric model

### v4

@@ IV Perturbative renormalization group results
For a fixed $g_{r}$ and an increasing $g_{i}$ with gradually enlarged non-Hermiticity, the system first flows to the strong-coupling phase for small $\theta$ and then flows to the weak-coupling phase for large $\theta$. Since the perturbative RG method is only expected to be valid and reliable for weak-coupling phase, it is intriguing to ask whether more exotic phases could occur in the strong or intermediate coupling regime. Notably, recent Bethe ansatz studies have clearly implied that there indeed emerges an extra phase in the intermediate region for intermediate $\theta$ [35], where additional degrees of freedom screen the local moment. This is termed as the Yu-Shiba-Rusinov (YSR) phase owing to its similarity with impurity states in superconductors [113, 114]. Since this emerging new phase is missed by the perturbative RG analysis, it is obvious that more careful non-perturbative approaches are indispensable to achieve a better and comprehensive understanding of the non-Hermitian Kondo effect.
V Bethe ansatz solution
@@
where $v_{f}$ denotes the Fermi velocity, $\psi_{k,\alpha}^{\dagger}(x),\psi_{k,\alpha}(x)$ are the creation and annihilation operators of conduction electrons (after the linearization of the spectrum around Fermi surface and the mapping to 1D model [115, 116, 117, 118, 119, 120, 121]) defined at coordinate $x$ with spin $\alpha$ and channel $k$. $J$ is the bare complex coupling. The corresponding Bethe ansatz equations are obtained as,
@@
The non-Hermitian Kondo model is characterized by two RG-invariant parameters [35], i.e., the Kondo energy scale $T_{K}$ and the non-Hermiticity parameter $\alpha$, given by,
@@
where $\nu=\frac{N}{L}$ is the density of electrons. For simplicity, we set $\widetilde{g}=1$ in the following calculations.
@@
where we have used the identity, $\int\text{sech}(\frac{\pi x}{n})dx=n$. The total spin $S=0$ implies that the impurity is fully screened.
@@
The spin excitations carrying spin 1/2 are termed as the spinons. Note that the spinon energy is complex, which consists of both a real part $E_{h}$ and an imaginary part $\Gamma_{h}$.
@@
which is a function of $\alpha$ and $\frac{E}{T_{K}}$. We then plot $\nu_{s}$ for different $\alpha$, as shown in Fig.2(a). When $0\leq\alpha\leq\pi/6$, the spinon DOS remains a pure Kondo behavior with a peak at $E=0$. When $\alpha\geq\pi/6$, the peak moves to $E/T_{K}=\sqrt{2\sin\alpha-1}$, until $\alpha$ approaches $\pi/2$. Meanwhile, the spinon DOS develops a delta peak, i.e., $\delta(E-T_{K})/2$ at $\alpha=\pi/2$. These results coherently indicate that the number of modes that screen the impurity decreases with $\alpha$, until reaching the case where the screening is contributed only by a single mode at $E/T_{K}=1$.
@@
The above Bethe ansatz solution is obtained for the general non-Hermitian multichannel Kondo model with isotropic Kondo exchange couplings, thus is not applicable for the $\mathcal{PT}$ symmetric case. For the $\mathcal{PT}$ symmetric case (with $J^{\prime}_{3}=J^{\prime}$, $J^{\prime}_{1}=J^{\prime\star}_{2}=J^{\prime}e^{i\theta}$), although we do not have access to the specific Bethe ansatz solution yet, we expect that only the contribution of the real part of impurity scattering could remain, since the imaginary part contributed by the conjugated $J^{\prime}_{1}$, $J^{\prime\star}_{2}$ pairs may cancel with each other, as will be implied by the NHNRG results below. Here, for the $\mathcal{PT}$ asymmetric case, we clearly see that the non-Hermitian effects can induce exotic phases absent in the Hermitian counterparts. Then, a urgent question to be answered is that how to obtain clear signatures of the emerging YSR phase, particularly regarding the thermodynamic quantities. We will further clarify this question in the next section using the NHNRG method.
VI non-Hermitian nrg approach
VI.1 Numerical implementations
In this section, we demonstrate the details of the implementation of NHNRG, focusing on its difference with conventional NRG for Hermitian systems. We note that a previous work by P. C. Burke and A. K. Mitchell [84] developed a NHNRG approach and solved the non-Hermitian single-channel quantum impurity model. Here, we further study the multichannel cases and focus on the thermodynamic quantities with varying the temperature (see Eq. (73) and (74) below).
The conventional NRG method involves several steps [127]. The first step is the logarithmical discretization. One reformulates the original impurity model using a one-dimensional (1D) energy representation for the bath electrons under a band width cutoff $\pm D$. The conduction band can be discretized to a series of energy intervals $D_{n}$ with a scaling parameter $\Lambda$, i.e.,
@@ VI.1 Numerical implementations
where $H_{N}$ is the first $N$ site Hamiltonian, and $h_{N+1}$ represents short-range terms containing the $(N+1)$-th site. In each iteration, we truncate the energy spectrum and dilate the kept eigenenergies, which gradually yields the low-energy physics. The success of this strategy is guaranteed by the energy scale separation of nearest sites in the Wilson chain, owing to the logarithmical discretization. The iteration diagonalization can then be stopped as long as a stable low-energy fixed point is reached. Such a diagonalization process also reveals information about the fixed points, as well as thermodynamic properties with varying energy scales.
@@
Although the two sets of wavefunctions are not orthogonal by themselves alone, they do form a biorthogonal basis that satisfies, [129]
@@
From above, one observes that the differences between Hermitian and non-Hermitian NRG lie in their different ways of diagonalization. For non-Hermitian systems, we need to calculate both left and right eigenvectors for each iteration step, as clarified in Eq.(57). This generates more detailed issues regarding the non-unitary transformation, truncation, and the degeneracy, etc, which will be explained in the following.
@@
When $\mathcal{PT}$ symmetry is not spontaneously broken, the energy spectrum will be strictly real. In this case, the metric $\eta$ is not only invertible but also positive. Thus, the thermodynamic quantities defined in Eq.(74) are completely valid and is analogous to those in conventional Hermitian systems [133, 129]. Whereas, when the $\mathcal{PT}$ symmetry is spontaneously broken, the eigenenergies will either be real or take place as complex conjugate pairs. Thus, the partition function $Z_{N}^{\rm LR}$ still maintains to be real, since we have $e^{-\beta E}+e^{-\beta E*}=2e^{-\beta{\rm Re}(E)}{\rm cos[Im}(\beta E)]$. In this case, it should be noted that the Boltzmann weight and the partition function is not positive-definite in general. Although this is forbidden for closed Hermitian systems, this feature reflects the essence of the non-Hermiticity, which physically originates from open systems with loss or gain. Moreover, for the model and parameters studied in this work, we will show in the following that the partition function remains positive for all energy scales, as indicated by Fig.6(b).
@@
Now we discuss more numerical details about solving the left and right eigenstates, which is complicated by degeneracies, as is also discussed comprehensively in Ref. [84]. The degeneracies often arise from symmetries, either at the Hamiltonian level or the emergent ones at low-energies. The problem can be partially overcome by decomposing the Hamiltonian into subspaces of different irreducible representations. Then, one only needs to solve certain reduced block matrices. Since the sizes of the blocks are much smaller than the full matrix, the computational cost is also reduced [134, 135]. Taking the $n=2$ NHMCK model as an example, one can utilize the ${\rm U}(1)_{\rm charge}\times{\rm SU}(2)_{\rm channel}$ symmetry. However, for symmetries emergent around the low-energy fixed points [136], the degeneracy cannot be avoided. For non-Hermitian matrices, most of numerical eigensolver packages do not support a biorthonormal output. One has to orthogonalize and normalize the eigenstates manually in each degenerate subspace. In fact, one can first solve the right eigenvectors with common eigensolvers, and then obtain the left eigenvectors by the inverse of the right one (when invertible), i.e.,
@@
Truncation is another key procedure in NRG calculations. In Hermitian cases, the usual scheme is just to drop out excited states higher than an energy cutoff $E_{\rm k}$, keeping the lowest $N_{\rm k}$ states. For non-Hermitian cases, the eigenenergies are generally complex and the cutoff needs a careful benchmark. In Ref. [84], several truncation schemes are discussed, and a “LowRe” scheme is demonstrated to yield the most stable and accurate numerical results given the cutoff $N_{\rm k}$. This scheme sorts the eigenvalues primarily by their real parts $\textrm{Re}(E)$ and secondarily by their imaginary parts $\textrm{Im}(E)$. We adopt the same cutoff scheme in our calculations in the main text, and provide a basic convergence analysis on different truncation schemes as well as the discretization parameter $\Lambda$ in Appendix. D.
VI.2 Non-Hermititian NRG results for the PT-asymmetric model
We now solve the NHMCK model taking $n=2$ (two-channel) as an example. Compared to models with more channels, it is of minimal computation cost and is already sufficient to demonstrate the key numerical features. The channel-symmetry is assumed since the proposed setup in Sec. III ensures exact symmetry between different channels. For the $\mathcal{PT}$-asymmetric model, we consider a general complex Kondo exchange coupling, $J=J_{0}e^{i\theta}$ with a tunable $\theta$ that quantifies the strength of non-Hermiticity. For the $\mathcal{PT}$-symmetric model, we assume anisotropic Kondo coupling that takes the form of ${\bf J}=(J_{x},J_{y},J_{z})=J_{0}(e^{i\theta},e^{-i\theta},1)$. Such an anisotropic model can also be realized by tuning the tunneling coefficient in the multi-junction setup, as discussed earlier. For $\theta=0$, both cases reduce to the conventional Hermitian two-channel Kondo model. In this section, we show the NRG results for the $\mathcal{PT}$-asymmetric model, while those for the $\mathcal{PT}$-symmetric model will be presented in the next section.
Figure 3: The energy spectrum for the $\mathcal{PT}$-broken $n=2$ NHMCK model obtained by non-Hermitian NRG for different $\theta$. The upper ((a)(c)(e)) and lower ((b)(d)(f)) panel show the real and imaginary part, i.e., Re(E)(E) and Im(E)(E), respectively. The maximal absolute value of the imaginary parts are marked for each plot in (b)(d)(f). The first 50 eigenenergies are plot here. The black curves show the calculated Re(E)(E) corresponding to the Hermitian two-channel Kondo model with $\theta=0$. For the $\mathcal{PT}$-asymmetric model, we take the numerical parameters as $\Lambda=6$, $N_{\rm k}=2048$, and adopt the “LowRe” truncation scheme. The same parameters are used in Fig.4.
@@ VI.2 Non-Hermititian NRG results for the PT-asymmetric model
The energy spectrum of the $\mathcal{PT}$-asymmetric model is shown in Fig.3. Notably, the spectrum deviates from the Hermitian result (the dark curves) as $\theta$ increases. For small $\theta$ , the real part of the spectrum Re(E)(E) agrees well with that of the Hermitian case (Fig.3(a)), with only a negligible imaginary part (Fig.3(b)). For intermediate $\theta$, Re(E)(E) displays obvious deviations at relatively high energy scales (Fig.3(c)), exhibiting larger imaginary part in the same energy range (Fig.3(d)). However, the system still flows to the Hermitian fixed point at low energies. For large $\theta$, the spectrum is completely distinct from the Hermitian, and the low-energy fixed point is also altered. The absolute values of the imaginary part Im($E$) also increase with $\theta$, as marked in Fig.3(b),(d),(f).
@@
Thermodynamic quantities for the $\mathcal{PT}$-asymmetric model also display interesting non-Hermitian effects. We calculate the impurity entropy $S_{\rm imp}$ and the impurity specific heat $C_{\rm imp}$ as a function of temperatures. They are defined as the thermodynamic quantities of the total system subtracting that of the bath [127],
$S_{\rm imp}(T)$ $=S_{\rm tot}(T)-S_{\rm bath}(T),$ (77)
$C_{\rm imp}(T)$ $=C_{\rm tot}(T)-C_{\rm bath}(T),$ (78)
where $S_{\mathrm{tot/bath}}(T)$ and $C_{\mathrm{tot/bath}}(T)$ are computed according to the LR-form defined above.
As shown in Fig.4(a), the model in fact exhibits three different regimes as $\theta$ increases. The first is the Kondo regime with weak non-Hermiticity ($\theta=0.10$) where $S_{\rm imp}$ monotonically evolves from ln2 to $\frac{1}{2}$ln2. This indicates the (over)screening process of the impurity with lowering temperature, similar to the Hermitian case. The second is the intermediate non-Hermiticity regime ($\theta=0.50$), where $S_{\rm imp}$ also evolves from ln2 to $\frac{1}{2}$ln2, but clearly in a non-monotonous behavior. The third is the strong non-Hermiticity regime ($\theta=1.00$), where the impurity entropy $S_{\rm imp}$ is close to ln 2 in both the UV and IR limit, displaying the local moment phase.
@@
We further compare the imaginary parts of the partition function $Z$ and the impurity entropy $S_{\rm imp}$ to their real parts in Fig.4(b) and (c). We find that the imaginary parts vanish as the NRG iteration goes to the IR fixed point, but can not be ignored at higher energies. Furthermore, the nonzero imaginary parts hallmark the characteristic energy scale where real parts of the impurity thermodynamic properties show YSR-like behaviors. The non-monotonic entropy and the overshoot of $\mathrm{ln}2$ indicate that there are extra DOFs with finite energy contributing to the impurity thermodynamics, which are totally induced by the non-Hermitian effects.
Figure 5: The NHNRG results of the energy spectrum for the $\mathcal{PT}$-symmetric $n=2$ NHMCK model for different $\theta$. For the $\mathcal{PT}$-symmetric model, we take the numerical parameters as $\Lambda=6$, $N_{\rm k}=2048$, and adopt the “LowRe” truncation scheme as well. The same parameters are used in Fig.6.
VII the pseudo-Hermitian model with PT symmetry
VII.1 Non-Hermitian NRG results for the PT-symmetric model